# Single-layer tensor network approach for three-dimensional quantum systems

### v1

@@ I Introduction
In this paper, we follow the simple update methodology to optimize the 3d iPEPS on the cubic lattice, whilst propose and test a single-layer computational procedure (based on ideas from Refs. [50, 51]) for calculating observables with it. The approach extensively utilizes the layered structure of the tensor network to reduce the scaling with the bond dimension at the expense of the enlargement of the elementary unit cell. The beneficial property of this method is its $D^{12}-D^{13}$ computational scaling (depending on specific details), which is comparable to the cost of the typical calculations with the 2d iPEPS approach. Let us emphasize that the latter corresponds to the cost of computation of observables and not the cost of the iPEPS optimization, which is typically smaller. Hence, this computation should be performed only once during the calculation (at least, if one follows the simple update strategy in the iPEPS optimization). Our approach is partially connected with a recent study [52] with the single-layer TN contraction applied to a classical 3d statistical mechanics problem.
@@ II Algorithm description
The optimization in terms of PEPO runs in several steps. First, we initialize the iPEPS wave function. We use the tensors with $D=1$ as the initial iPEPS, since this initialization speeds up the convergence according to our observations. Second, we determine the PEPO, which approximates the operator $\exp(-Hdt)$. This can be done, e.g., by using the $W^{I}$ approach or other cluster-based methods [54, 55, 56]. After that, we repeatedly apply the PEPO to the iPEPS wave function, as in Fig. 1(b). The bond dimension $D$ grows with every PEPO application. Hence, after several first applications one needs to start truncating the iPEPS bond dimension back to some fixed target bond dimension $D$. The truncation can be performed by using the superorthogonal canonical form [57, 58, 59, 60], which is defined by the condition shown in Fig. 2(b). To truncate the iPEPS bond dimension, we iteratively transform it to the canonical form, as in Fig. 2(c), and then truncate the bonds according to the magnitudes of diagonal elements of $\lambda$ (we truncate the smallest diagonal elements). Note that in the superorthogonal canonical form the matrices $\lambda$ are diagonal and positive.
@@ II.3 Boundary iPEPS: Simple update scheme
Upon completing the iPEPS optimization, either by applying the above PEPO scheme or Trotterized gates, it is necessary to compute observables with the obtained wave function. To this end, we need to find an efficient method to compute the wave function norm $\langle\Psi|\Psi\rangle$ and the local operator insertions $\langle\Psi|\hat{O}|\Psi\rangle$, where $\hat{O}$ is the local operator. The wave function norm can be cast in the form of contraction of the two infinite 3d tensor networks, which consists of the double-layer tensors $A=T^{p}_{lrudio}\bar{T}^{p}_{l_{2}r_{2}u_{2}d_{2}i_{2}o_{2}}$, where the pairs of indices, e.g., $l$ and $l_{2}$, are combined into a double-layer index, as shown in Fig. 3. This double-layer index has the dimension $D^{2}$ (note that we absorb the bond matrices into the tensors $T$).
Figure 3: (a) Definition of the double-layer tensor $A$ and (b) infinite 3d tensor network composed of the these tensors (the transfer matrix $M$ consists of all tensors $A$ on the horizontal plane); (c) The left boundary PEPS (bPEPS) consists of identical tensors $B$ and positive matrices $\lambda$ placed on its virtual links. (d) Each tensor $B$ has the bond dimension $\chi_{b}$ for its four virtual indices.
@@
To determine the leading eigenvectors of $M$, we follow the suggestion from Ref. [44], i.e., approximate these leading eigenvectors as the two-dimensional iPEPS of the bond dimension $\chi_{b}$ with the role of physical index replaced by the combined index of double-layer tensor $A$ in the transverse direction to the transfer matrix plane. For example, if the transfer-matrix layers are in the $xy$ plane, then the indices $i,o$ are transverse, and physical indices of the left iPEPS tensors are $i,i_{2}$, while the physical indices of right boundary iPEPS are $o,o_{2}$. In total, the left boundary iPEPS consists of the rank-6 tensors $B^{ii_{2}}_{lrud}$, with $i$ and $i_{2}$ of the dimension $D$ (from the $T$ tensors of 3d iPEPS), while $l,r,u,d$ are the auxiliary indices of the dimension $\chi_{b}$, as shown in Fig. 3(d). Note that in the case of the enlarged unit cell of the bulk 3d iPEPS, the boundary iPEPS tensors $B$ must be also taken with the enlarged unit cell, which is the projection of 3d unit cell on the boundary plane.
Our nest step is to determine the left boundary iPEPS tensors $B$. In this section, we follow another suggestion from Ref. [44] and adapt the simple update scheme to find $B$. Still, our implementation of the this update is different from Ref. [44]. We propose to use explicitly the double-layer structure of the transfer matrix tensors $A$ to reduce the scaling of the simple update calculations to $\chi_{b}^{5}D^{7}$.
@@
Figure 4: (a,b) Transfer matrix $M$ can be applied to the bPEPS in a layered fashion. First, the tensors $T$ are absorbed into the bPEPS tensors $B$ to obtain the new enlarged tensor $B^{\prime}$. Next, the bond dimension of the tensor $B^{\prime}$ is truncated back to the original dimension $\chi_{b}$, which results in the auxiliary (intermediate) tensor $B_{a}$. (c,d) Similar procedure with the second layer composed of $T^{\dagger}$ tensors: We absorb $T^{\dagger}$ into $B_{a}$ and then truncate the enlarged tensors $B^{\prime}_{a}$ back to the original dimension $\chi_{b}$. This yields the original bPEPS tensors $B$. The iterations are repeated till convergence of the tensor $B$ and the corresponding matrix $\lambda$.
@@ II.4 Boundary CTMRG: Three-layer approach
Figure 5: (a) The infinite 2d tensor network necessary for the local observables computation. (b) The bond matrices $\lambda$ can be absorbed into the bPEPS tensors $B$, resulting in the new boundary tensors $b$. The boundary tensors $b$ can sandwich the bulk tensor $A$ into the new single layer tensor $t$ of the bond dimension $D^{2}\chi_{b}^{2}$. (c) Alternatively, one can absorb the tensors $T$ (from $A$) into the two bPEPS layers with the tensors $B^{\prime}$. The structure can be further approximated (by the leg dimension truncation back to $\chi_{b}$) to the 2-layer tensor network of $B_{a}$.
@@
Figure 8: The illustration of the mappings of the 4-layer tensor network to the single-layer networks with $3\times 3$ or $4\times 4$ unit cells.
@@ III Results for the Heisenberg model
where we set $J=1$ (antiferromagnetic coupling), ${\bf S}_{i}$ are the local spin operators with the conventional relations to the Pauli matrices ($S^{\alpha}=\sigma^{\alpha}/2$ with $\alpha=X,Y,Z$), and $\langle ij\rangle$ runs over nearest-neighbor pairs of sites $i$ and $j$. The ground state of the model corresponds to the gapless antiferromagnetic phase with a two-site unit cell (containing sites $a$ and $b$). Still, it can mapped to the system with a single-site unit cell by using the unitary transformation $R_{a}$ on all the sites of one sublattice ($i\in a$). We choose the wave function to be real, so that the magnetization lies in the $XZ$ plane, and then perform the unitary transformation $R_{a}=\exp(i\pi S^{Y})=i\sigma^{Y}$, which rotates the magnetization on each site $i\in a$ by $\pi$ angle around the $Y$ axis. After this unitary transformation, the bulk iPEPS becomes completely homogeneous with identical tensors on all the sites. Note also that the wave function preserves lattice rotational and reflection symmetries. We have used the ITensors numerical package [66] in our calculations.
@@
Let us now focus on the single-layer calculations and determine the characteristic values of $\chi_{b}$ and $\chi$, which yield reliable results. In Fig. 9(b) we show the convergence of magnetization with $\chi$ for $D=7$ and $\chi_{b}=16$ (these were the maximal $D$ and $\chi_{b}$ we were able to reach). We see that the convergence is reached at $\chi\approx 200$. In general, throughout our analysis we observed an approximate relation $\chi\approx\chi_{b}^{2}$ for the final convergence in $\chi$.
@@
Our next goal is to collect the results for different bond dimensions $D$ and to extrapolate them to the infinite-$D$ limit. The correlation length $\xi$ is typically viewed as a reliable measure to extrapolate the calculated observables. Note that in the single-layer approach the correlation length can be obtained at the same computational cost as in 2d problems. For the extrapolation we employ the following scaling relations [67]:
@@
The results of the fit are shown in Fig. 11. In the limit $\xi\to\infty$ we obtain $E=-0.90237(2)$ for the energy and $m^{2}=0.1781(3)$ for the magnetization per site. The fitting of the data is performed from the iPEPS simulations with $D\in[4,7]$ (the results for $D=3$ appear not completely consistent with the fitting ansatz). Let us additionally note that the correlation length is calculated for the largest available $\chi_{b}$ and $\chi$ at the given $D$ and is not additionally extrapolated to the limits $\chi\to\infty$ and $\chi_{b}\to\infty$. Hence, it could explain possible minor errors in the estimated correlation length values. Nevertheless, our estimates agree reletively well with the available published results from the quantum Monte-Carlo approach, $E_{\rm QMC}=-0.902325(11)$ and $m^{2}_{\rm QMC}=0.1786(4)$ [44].
@@ IV Conclusions
In this study, we performed the tensor network calculations for the 3d Heisenberg model on the cubic lattice. We intensively used the multilayer structure of the tensor network for the calculation of observables in order to diminish the computational cost both during the boundary iPEPS search and during the final CTMRG contraction. As a result, we were able to perform calculations up to $D=7$ without usage of U(1) symmetries.
There are several research directions for the future studies. In particular, introducing U(1) symmetries [68, 69] may allow reaching even higher bond dimensions $D$. Additional research directions concern the algorithm development for larger unit cells, as well as its application to other lattice geometries and fermionic models [70, 71, 72]. Furthermore, it would be also interesting to apply the full update optimization [73, 57] of 3d tensor networks by means of the proposed single-layer CTMRG environments. The latter looks manageable to us, at least, for $D=3$ (and possibly for $D=4$). It may be also possible to generalize the multilayer approach of observables calculation to finite temperature quantum systems [37, 38] and to classical 3d statistical mechanics models [74, 75, 76].
@@ Acknowledgements.
The authors acknowledge support by the National Research Foundation of Ukraine under the call “Excellent science in Ukraine” (2024–2026).

### v2

@@ I Introduction
In this paper, we follow the simple update methodology to optimize the 3d iPEPS on the cubic lattice, whilst proposing and testing a single-layer computational procedure (based on ideas from Refs. [50, 51]) for calculating observables with it. The approach extensively utilizes the layered structure of the tensor network to reduce the scaling with the bond dimension at the expense of the enlargement of the elementary unit cell. The beneficial property of this method is its $D^{12}-D^{13}$ computational scaling (depending on specific details), which is comparable to the cost of the typical calculations with the 2d iPEPS approach. Let us emphasize that the latter corresponds to the cost of computation of observables and not the cost of the iPEPS optimization, which is typically smaller. Hence, this computation should be performed only once during the calculation (at least, if one follows the simple update strategy in the iPEPS optimization). Our approach is partially connected with a recent study [52] with the single-layer TN contraction applied to a classical 3d statistical mechanics problem.
@@ II Algorithm description
The optimization in terms of PEPO runs in several steps. First, we initialize the iPEPS wave function. We use the tensors with $D=1$ as the initial iPEPS, since this initialization speeds up the convergence according to our observations. Second, we determine the PEPO, which approximates the operator $\exp(-Hdt)$. This can be done, e.g., by using the $W^{I}$ approach or other cluster-based methods [54, 55, 56] (see also Appendix  A for more details). After that, we repeatedly apply the PEPO to the iPEPS wave function, as illustrated in Figs. 1(b), 1(c), and 2(a). The bond dimension $D$ grows with every PEPO application. Hence, after several first applications one needs to start truncating the iPEPS bond dimension back to some fixed target bond dimension $D$. The truncation can be performed by using the superorthogonal canonical form [57, 58, 59, 60], which is defined by the condition shown in Fig. 2(b). This condition should hold for all tensors and bonds of the iPEPS wave function and it is a natural generalization of the canonical form of matrix product states.
To truncate the iPEPS bond dimension, we iteratively transform it to the canonical form, as in Fig. 2(c), and then truncate the bonds according to the magnitudes of diagonal elements of $\Lambda$ (we truncate the smallest diagonal elements). Note that in the superorthogonal canonical form, the matrices $\Lambda$ are diagonal and positive. The transformation into the canonical form can be performed as shown in Fig. 2(c): First, one can take one particular bond of the iPEPS wave function with the two tensors $T$ on this bond and absorb all matrices $\Lambda$ on the adjacent bonds into these tensors. Next, one can decompose these tensors by means of the QR decomposition, as we also show in Fig. 2(c). The resulting matrices $R_{L}$ and $R_{R}$ can be viewed as obstructions to the conditions in Fig. 2(b). To get rid of these matrices, one can simultaneously insert the matrices $R_{L}$ and $R_{R}$ and their inverses on the bonds. Inverses can be then absorbed into the tensors $T$, while the matrices $R_{L}$ and $R_{R}$ are absorbed into the bond matrix $\Lambda$. The resulting bond matrix (which is no longer diagonal) is now decomposed with the singular value decomposition (SVD), $R_{L}\Lambda R_{R}=U\Lambda^{\prime}V$, and the corresponding SVD unitary matrices $U$ and $V$ are also absorbed into the tensors $T$, while the SVD spectra form the new positive and diagonal matrices $\Lambda^{\prime}$. The iterations can be repeated until convergence of the bond matrices $\Lambda$.
@@ II.3 Boundary iPEPS: Simple update scheme
Upon completing the iPEPS optimization, either by applying the above PEPO scheme or Trotterized gates, it is necessary to compute observables with the obtained wave function. To this end, we need to find an efficient method to compute the wave function norm $\langle\Psi|\Psi\rangle$ and the local operator insertions $\langle\Psi|\hat{O}|\Psi\rangle$, where $\hat{O}$ is the local operator. The wave function norm can be cast in the form of contraction of the two infinite 3d tensor networks, which consists of the double-layer tensors $A=T^{p}_{lrudio}\bar{T}^{p}_{l_{2}r_{2}u_{2}d_{2}i_{2}o_{2}}$, where the pairs of indices, e.g., $l$ and $l_{2}$, are combined into a double-layer index, as shown in Fig. 3(a). This double-layer index has the dimension $D^{2}$ (note that we absorb the bond matrices into the tensors $T$).
Figure 3: (a) Definition of the double-layer tensor $A$ and (b) infinite 3d tensor network composed of these tensors (the transfer matrix $M$ consists of all tensors $A$ on the horizontal plane); (c) The left boundary PEPS (bPEPS) consists of identical tensors $B$ and positive matrices $\lambda$ placed on its virtual links. (d) Each tensor $B$ has the bond dimension $\chi_{b}$ for its four virtual indices.
@@
To determine the leading eigenvectors of $M$, we follow the suggestion from Ref. [44], i.e., approximate these leading eigenvectors as the two-dimensional iPEPS of the bond dimension $\chi_{b}$ with the role of physical index replaced by the combined index of double-layer tensor $A$ in the transverse direction to the transfer matrix plane, as depicted in Fig. 3(c). For example, if the transfer-matrix layers are in the $xy$ plane, then the indices $i,o$ are transverse, and physical indices of the left iPEPS tensors are $i,i_{2}$, while the physical indices of right boundary iPEPS are $o,o_{2}$. In total, the left boundary iPEPS consists of the rank-6 tensors $B^{ii_{2}}_{lrud}$, with $i$ and $i_{2}$ of the dimension $D$ (from the $T$ tensors of 3d iPEPS), while $l,r,u,d$ are the auxiliary indices of the dimension $\chi_{b}$, as shown in Fig. 3(d). Note that in the case of the enlarged unit cell of the bulk 3d iPEPS, the boundary iPEPS tensors $B$ must be also taken with the enlarged unit cell, which is the projection of the 3d unit cell on the boundary plane.
Our next step is to determine the left boundary iPEPS tensors $B$. In this section, we follow another suggestion from Ref. [44] and adapt the simple update scheme to find $B$. Still, our implementation of this update is different from Ref. [44]. We propose to use explicitly the double-layer structure of the transfer matrix tensors $A$ to reduce the scaling of the simple update calculations to $\chi_{b}^{5}D^{7}$.
@@
Figure 4: (a) and (b) Transfer matrix $M$ can be applied to the bPEPS in a layered fashion. First, the tensors $T$ are absorbed into the bPEPS tensors $B$ to obtain the new enlarged tensor $B^{\prime}$. Next, the bond dimension of the tensor $B^{\prime}$ is truncated back to the original dimension $\chi_{b}$, which results in the auxiliary (intermediate) tensor $B_{a}$. (c) and (d) Similar procedure with the second layer composed of $T^{\dagger}$ tensors: We absorb $T^{\dagger}$ into $B_{a}$ and then truncate the enlarged tensors $B^{\prime}_{a}$ back to the original dimension $\chi_{b}$. This yields the original bPEPS tensors $B$. The iterations are repeated till convergence of the tensor $B$ and the corresponding matrix $\lambda$.
@@ II.4 Boundary CTMRG: Three-layer approach
Figure 5: (a) The infinite 2d tensor network necessary for the local observables computation. (b) The bond matrices $\lambda$ can be absorbed into the bPEPS tensors $B$, resulting in the new boundary tensors $b$. The boundary tensors $b$ can sandwich the bulk tensor $A$ into the new single layer tensor $t$ of the bond dimension $D^{2}\chi_{b}^{2}$. (c) Alternatively, one can absorb the tensors $T$ (from $A$) into the two bPEPS layers with the tensors $B^{\prime}$. The structure can be further approximated (by the leg dimension truncation back to $\chi_{b}$) to the two-layer tensor network of $B_{a}$.
@@
We should note that the introduction of enlarged unit cells leads to the constant overhead in the computational time and memory requirements (e.g., these increase in 16 times for the $4\times 4$ unit cell). Still, in practice, the differences in scalings with the bond dimension become more crucial than these constant factors. In addition, the developed approach has the full potential to be generalized to the iPEPS wave functions with the larger original unit cells. For example, the $4\times 4$ mapping of the original unit cell of the size $l\times m$ (after projection on the plane) results in the single-layer tensor network with the unit cell of the size $4l\times 4m$ (see Ref. [51] for 2d case).
The converged CTMRG environments can also be used to obtain the reduced density matrices of the 3d iPEPS wave functions, which can be further employed in the local observables computation. For this purpose, one should leave open physical indices of the bulk tensor $T$ on one particular site and replace all other sites (and boundary iPEPS tensors) with the CTMRG environments. The separation of the tensors $T$ and $T^{\dagger}$ to different sites of the enlarged unit cell and the combination of physical indices with the virtual indices may lead to potential asymmetry between the wave-function tensors and their conjugates. This may become an additional source of the absence of hermiticity or positivity in the resulting reduced density matrices. Still, in practice, these non-Hermitian parts of the density matrices converge to zero with an increase of the CTMRG bond dimension $\chi$.
Figure 8: The illustration of the mappings of the four-layer tensor network to the single-layer networks with $3\times 3$ or $4\times 4$ unit cells.
@@ III Results for the Heisenberg model
where we set $J=1$ (antiferromagnetic coupling), ${\bf S}_{i}$ are the local spin operators with the conventional relations to the Pauli matrices ($S^{\alpha}=\sigma^{\alpha}/2$ with $\alpha=X,Y,Z$), and $\langle ij\rangle$ runs over nearest-neighbor pairs of sites $i$ and $j$. The ground state of the model corresponds to the gapless antiferromagnetic phase with a two-site unit cell (containing sites $a$ and $b$). Still, it can be mapped to the system with a single-site unit cell by using the unitary transformation $R_{a}$ on all the sites of one sublattice ($i\in a$). We choose the wave function to be real, so that the magnetization lies in the $XZ$ plane, and then perform the unitary transformation $R_{a}=\exp(i\pi S^{Y})=i\sigma^{Y}$, which rotates the magnetization on each site $i\in a$ by $\pi$ angle around the $Y$ axis (this ansatz is an example of more general spiral PEPS approach [66], which is also capable of describing incommensurate phases). After this unitary transformation, the bulk iPEPS becomes completely homogeneous with identical tensors on all the sites. Note also that the wave function preserves lattice rotational and reflection symmetries. We have used the ITensors numerical package [67] in our calculations.
@@
Let us now focus on the single-layer calculations and determine the characteristic values of $\chi_{b}$ and $\chi$, which yield reliable results. In Fig. 9(b) we show the convergence of magnetization with $\chi$ for $D=7$ and $\chi_{b}=16$ (these were the maximal $D$ and $\chi_{b}$ we were able to reach). We see that the convergence is reached at $\chi\approx 200$. In general, throughout our analysis, we observed an approximate relation $\chi\approx\chi_{b}^{2}$ for the final convergence in $\chi$.
@@
Our next goal is to collect the results for different bond dimensions $D$ and to extrapolate them to the infinite-$D$ limit. The correlation length $\xi$ is typically viewed as a reliable measure to extrapolate the calculated observables. Note that in the single-layer approach, the correlation length can be obtained at the same computational cost as in 2d problems. For this purpose, we employ the scaling relations from Ref. [68] (while the procedure for extrapolation of iPEPS results with the correlation-length scaling was proposed in Ref. [69]):
@@
The results of the fit are shown in Fig. 11. In the limit $\xi\to\infty$ we obtain $E=-0.90237(2)$ for the energy and $m^{2}=0.1781(3)$ for the magnetization per site. The fitting of the data is performed from the iPEPS simulations with $D\in[4,7]$ (our results for $D=3$ with $\xi=0.53$, $E(\xi)=-0.89960$, and $m^{2}(\xi)=0.1935$, which are obtained at $\chi_{b}=8$, appear not completely consistent with the fitting ansatz and not shown in Fig. 11 for the sake of visibility of the main data). Let us additionally note that the correlation length is calculated for the largest available $\chi_{b}$ and $\chi$ at the given $D$ and is not additionally extrapolated to the limits $\chi\to\infty$ and $\chi_{b}\to\infty$. Hence, it could explain possible minor errors in the estimated correlation length values. Nevertheless, our estimates agree relatively well with the available published results from the quantum Monte-Carlo approach, $E_{\rm QMC}=-0.902325(11)$ and $m^{2}_{\rm QMC}=0.1786(4)$ [44]. One can also compare our results with the extrapolated iPEPS values from Ref. [44]: $m^{2}=0.1826(2)$ and $E=-0.9024(1)$. We see that the energy estimates agree well with those from Ref. [44], while the magnetization is significantly lower indicating that the estimates are closer to the QMC prediction (in fact, at $D=6$ our data for $m^{2}$ practically approach the extrapolated iPEPS result from Ref. [44]).
@@ IV Conclusions
In this study, we performed the tensor network calculations for the 3d Heisenberg model on the cubic lattice. We intensively used the multilayer structure of the tensor network for the calculation of observables in order to diminish the computational cost both during the boundary iPEPS search and during the final CTMRG contraction. As a result, we were able to perform calculations up to $D=7$ without the usage of U(1) symmetries.
There are several research directions for the future studies. In particular, introducing U(1) symmetries [70, 71] may allow reaching even higher bond dimensions $D$. Additional research directions concern the algorithm development for larger unit cells, as well as its application to other lattice geometries and fermionic models [72, 73, 74]. Furthermore, it would be also interesting to apply the full update optimization [75, 57] of 3d tensor networks by means of the proposed single-layer CTMRG environments. The latter looks manageable to us, at least, for $D=3$ (and possibly for $D=4$). It may be also possible to generalize the multilayer approach of observables calculation to finite temperature quantum systems [37, 38] and to classical 3d statistical mechanics models [76, 77, 78]. Another relevant research direction is a more comprehensive analysis of the scalings of the CTMRG and boundary iPEPS bond dimensions with $D$, since the results reported in this study should be considered as particular observations valid for the specific 3d Heisenberg model.
@@ Acknowledgements.
The authors acknowledge support by the National Research Foundation of Ukraine under the call “Excellent Science in Ukraine”, project No. 2023.03/0073 (2024–2026).
Appendix A $W^{I}$ PEPO construction
Let us briefly discuss how to generalize the construction of the matrix product operator (MPO) of the type $W^{I}$ for the evolution (time-stepper) operator $\exp(-Hdt)$ [54] to 3d PEPO. The main idea of the $W^{I}$ MPO approach is to construct an operator, which properly captures all terms in the expansion of $\exp(-Hdt)$, which are of the first order in $dt$ and all terms of the higher orders in $dt$ with the condition that the support of individual operators in the expansion does not overlap. For example, let us take the Hamiltonian $H=\sum_{\langle ij\rangle}h_{ij}$, where the sum runs over all nearest neighbor pairs of sites. Then, this construction captures all terms of the type $dt^{2}h_{ij}h_{kl}$ if all the sites $i,j,k,l$ are different (i.e., the support of operators $h_{ij}$ and $h_{kl}$ does not overlap). Hence, the natural generalization of the $W^{I}$ method to PEPO on arbitrary lattices should also capture all these operators with non-overlapping support.
We are interested in the $W^{I}$ construction for the Heisenberg model on the cubic lattice. First, note that, since our iPEPS is constructed by means of additional unitary transformations $i\sigma^{Y}$ on one sublattice (see Sec. III), the effective Hamiltonian on the bond with account of these rotations becomes $h_{kl}=-S_{k}^{X}S_{l}^{X}-S_{k}^{Z}S_{l}^{Z}-(iS_{k}^{Y})(iS_{l}^{Y})$. In Fig. 12 we show the non-zero PEPO terms for the Heisenberg model (there we show only non-zero terms for the $x$ axis of the cubic lattice; the non-zero terms for other axes are constructed analogously).
Figure 12: Construction of the $W^{I}$ PEPO for the Heisenberg model. We fix virtual indices to the values $0,1,2,3$ and obtain operators acting on the physical indices. Note that here we show only terms corresponding to operators acting on the bonds along the $x$ axis of the cubic lattice. The similar terms corresponding to $y$ and $z$ axes have the same structure. All other terms (e.g., with several indices equal to $1$) vanish.
Within this approach, the PEPO has a virtual bond dimension $4$. If all virtual indices are set to zero, PEPO acts on the given site as the identity operator. Next, e.g., if the right virtual index is equal to one, then the PEPO site to the right must have its left index to be equal to one, which results in the operator $dtS_{i}^{X}S_{j}^{X}$ (if the index is equal to two or three, this results in $-dtS_{i}^{Y}S_{j}^{Y}$ or $dtS_{i}^{Z}S_{j}^{Z}$, respectively). Since all other virtual indices for these two sites are zero, the operators on other sites cannot overlap with the operator $S_{i}^{X}S_{j}^{X}$. However, all zero virtual indices also mean that the other sites are completely indifferent to the presence of the operator $S_{i}^{X}S_{j}^{X}$. Hence, all types of operators $h_{ij}h_{kl}$ appear in the PEPO expansion as long as the sites $i,j,k,l$ are different.
This PEPO construction has several important properties: it is real-valued and it is symmetric under reflections and rotations. Hence, the PEPO application to 3d iPEPS wave function does not spoil its properties and these can be preserved in the process of optimization. There are several possible generalizations of this construction to other systems. In particular, it can be modified to capture interactions beyond nearest neighbors or even exponentially decaying long-range interactions. Other methods from Ref. [54], in particular, the $W^{II}$ approach can also be generalized to different lattices, but the precise recipe of this generalization is beyond the scope of the current study.